On September 30, 2021, Bricker Enterprises purchased a machine for $200,000. The estimated service life is 10 years with a $20,000 residual value. Bricker records partial-year depreciation based on the number of months in service.

Depreciation for 2022, using the double-declining-balance method, would be:

Multiple Choice

  • $32,000.

  • $34,000.

  • $38,000.

  • $40,000.
Answer #1

Depreciation rate = 200/useful life = 200/10 = 20%

Depreciation for 2021 = 200,000 * 20% * 3/12 = 10,000

Depreciation for 2022

= (200,000-10,000)*20%

= 38,000

Option C is the answer